Sales (40,000 units @ Br. 25 each) ............                   Br. 100,000

Less: Variable Costs..............................                               70,000

Contribution Margin .............................                             30,000

Less: Fixed Costs .................................                              18,000

Net Profit ..........................................                                       12.000

  1. The sales needed to earn a profit of 20% on sales
1
Expert's answer
2022-01-28T08:13:45-0500

Solution:

Target profit=sales + 20%=100,000+(100,000*20%)=120,000


Target sales= (Fixed costs + target profit) /contribution margin ratio


Contribution margin ratio =contribution margin/sales =30,000/100,000=0.3

Target sales =(18,000+120,000)/0.3= Br. 460,000


Target sales needed=460,000